A rescue plane wants to drop supplies to isolated mountain climbers on a rocky ridge 235m blelow the plane. If the pilot is moving at 250 KPH, how far must he drop the supplies?
1
Expert's answer
2018-10-01T13:58:08-0400
v=250/3.6 m/s.
We need to find the time of flight:
h=(gt^2)/2→t=√(2h/g)
the distance will be:
d=vt=v√(2h/g)
d=250/3.6 √((2(235))/9.8)=480 m.